# Federico Echenique and Farzad Pourbabaee

### v2

@@ 3.3 Prior Beliefs and Welfare-Improving Trade
Our third result concerns agents with multiple priors, and the size of the sets of prior beliefs that they may posses. In Section 3.1, we quantified the upper contour sets of individual agents and their sums. Here we instead follow Yaari (1969) to interpret a vector that supports the upper contour sets, at some contingent consumption, as a prior belief. We depend on a relaxation of the equivalence between the existence of a common prior and Pareto efficiency of an allocation. This equivalence has been studied in a number of previous works (e.g., Billot et al., 2000; Ng, 2003; Rigotti et al., 2008; Gilboa et al., 2014; Ghirardato and Siniscalchi, 2018). We shall follow Rigotti et al. (2008) quite closely here.
We consider an exchange economy $\mathcal{E}$ with no aggregate uncertainty. Importantly, here on we will not require convexity of preferences. The aggregate endowment is the same across all states of the world: $\omega=(\bar{\omega},\ldots,\bar{\omega})$. We quantify the space of all allocations, denoted by $\mathcal{F}_{\bar{\omega}}$, by the magnitude
$\rho\coloneqq 2\bar{\omega}^{-1}\max_{f\in\mathcal{F}_{\bar{\omega}}}\sum_{i\in I}\lVert f_{i}\rVert\,.$ (3.7)
@@
The set $\mathcal{B}_{i}(f)$ is a convex and closed (hence compact) subset of the $d$-dimensional probability simplex $\Delta S$. One may also interpret a vector in $\mathcal{B}_{i}(f)$ as the set of normalized prices that support the consumption of the act $f$ [^7] The price interpretation is suitable for the exercise in Section 3.1, where we could have phrased our results in terms of the Walrasian theory of general equilibrium when there is no uncertainty and $S$ captures the set of available consumption goods. Here we are more tightly following the story of uncertain consumption in financial markets.
@@
$\mathcal{B}_{i}(f_{i})^{\delta}\coloneqq\{\nu\in\Delta S:\inf_{\mu\in\mathcal{B}_{i}(f_{i})}\lVert\nu-\mu\rVert<\delta\}\,.$
@@
Leveraging this result, in the following theorem we examine the volume of the prior sets as the dimension $d$ grows large. Before we state the result, we need to introduce some further notation. For a subset $J\subseteq I$, denote its complement by $J^{c}$, and define $\mathcal{B}_{J}(f_{J})\coloneqq\bigcap_{j\in J}\mathcal{B}_{j}(f_{j})$. We often drop the allocation $f$ from the argument of subjective belief sets, when it is understood from the context.
@@
Let $\mathcal{E}\colon I\to\mathcal{P}\times\mathsf{E}^{d}_{+}$ be an exchange economy with no aggregate uncertainty.
If the allocation $f$ is $\varepsilon$-Pareto dominated, then there exists a constant $c>0$, such that for every proper subset $J\subset I$,
@@
The message of the theorem is clearest in a two-agent economy. The existence of a $\varepsilon$-Pareto improving trade is then only possible when one of the two agents has a small set of subjective beliefs, as measured through its Euclidean volume. We expand on this interpretation in Section 4.1.
More generally, in an economy with many agents, Theorem 3 says that, no matter how we partition the set of agents in two, at least one of the groups of agents must have diverse set of priors. The volume of the intersection of their sets of priors must be small. If we think of the intersection of the set of priors as corresponding to a representative agent for the group (a notion that may be formalized), then this agent must be close to being ambiguity neutral: see again the discussion in Section 4.1 for a sufficient condition under which the small volume conclusion leads to a behavioral property that bounds the degree of ambiguity aversion.
@@ 4.1 Behavioral Implications of Theorem 3
In this section we study the behavioral implications of Theorem 3. For the purposes of this discussion, we consider a setting with two agents. The message of our theorem is that, if the current risk-sharing arrangement between the two agents is strongly Pareto dominated (namely if there exists a possibility of a strongly welfare improving ex-ante trade), then the volume of the underlying belief set of at least one agent must be very small. The smaller is this volume, the closer is that agent to being ambiguity neutral — a connection we seek to quantify in this section. [^8] For an overview of ambiguity and ambiguity aversion, see Machina and Siniscalchi (2014) and Gilboa and Marinacci (2016).
@@
We further assume that $\Pi$ has constant width, namely that the distance between any distinct parallel supporting hyperplanes of $\Pi$ (residing on the $d$ dimensional probability simplex) is constant. This assumption is crucial for our use of the result of Schramm (1988), but one can imagine a quantitative relaxation of the assumption of constant width and that it would translate into a modified bound on the degree of ambiguity aversion.
@@
$\theta(f)\coloneqq\max_{\mu\in\Pi}f\cdot\mu-\min_{\mu\in\Pi}f\cdot\mu\,.$
When $\Pi$ has constant width, $\theta(f)$ becomes constant in $f$, and we write $\theta(f)\equiv\theta$. Thus we can take $\theta$ as a measure of ambiguity aversion. Its meaning is the difference in how an ambiguity averse and ambiguity loving agent would evaluate any given normalized act. A small degree of ambiguity aversion corresponds to a small difference between the ambiguity averse and ambiguity loving evaluation of an act.
@@
Under the conditions of Theorem 3, let $\Pi$ coincide with the set of priors with the smaller volume, and suppose that it has constant width $\theta$. Then there exists a universal constant $c>0$ such that
@@
4.2 On the Magnitude of $\varepsilon$
The interpretation of our results depends on the magnitude of two numbers: the number of states $d$, and the fraction $\varepsilon$ accounted for in a utility improvement. When $d$ is large while $\varepsilon$ remains constant, the volumes discussed in Theorems 1 and 2 shrink to zero. So one could question our interpretation by arguing that we should use a small value of $\varepsilon$ when the number of states $d$ is large. In particular, one might argue that we should impose that $\varepsilon=O(1/\sqrt{d})$. We would disagree, however, essentially because we think of $\varepsilon$ as a dimension-free fraction.
@@ 4.2 On the Magnitude of $\varepsilon$
Second, taking $\varepsilon=O(1/\sqrt{d})$ is problematic because it seems very hard to reconcile with the common practice of using a numerical objective function in calculating approximately optimal outcomes. As we discussed above, many applications make use of a homothetic preference and a resulting utility function that is homogenous. Examples include the max-min representation in choice under uncertainty, or the Cobb-Douglas utility in consumer choice. In this case, the $\varepsilon$ equals that tolerance level assumed in the agents’ maximization problem: the number $\varepsilon$ is then measured in “utils,” the same unit of account as used for the utility function. Utils are, however, dimension free. In a model with many states of the world, we would hardly be allowing for any relaxation in our notion of approximate optimality.
Finally, we should emphasize the reinterpretation of our results using the coefficient of resource utilization (see Corollary 2). The CRU is usually thought of as a fraction of national income, also a dimension free measure, and one that one might expect is constant, or even grow, with the size of the economy.
@@ 5 Methodology and Proofs
At a high level, the primary concept underpinning our proofs is that in high-dimensional metric spaces, and under relatively mild conditions, probability measures tend to concentrate. As a result, for a probability measure $\mu$ and a subset $A$ that encompasses at least half of the probability space, namely $\mu(A)\geq 1/2$, the metric extension $A^{\delta}=\{z:\mathsf{dist}(z,A)<\delta\}$ covers a substantial portion and growing of the unit measure. In particular, the complement $1-\mu(A^{\delta})$ diminishes rapidily, often exhibiting exponential decay with respect to the dimension $d$. These types of concentration bounds are commonly referred to by Isoperimetric inequalities. Their importance lie on the independence of the concentration rate from the set $A$. The source of our results, that certain welfare improvements have a vanishingly small probability, independently of the shape of the agents’ preferences, can be traced to basic results in the theory of Isoperimetric inequalities and concentration of measure.
A consequence of the uniform concentration of measure is that, if two subsets $A$ and $B$ are separated with a positive distance, then as the dimension $d$ grows, the measure of at least one of them must be exponentially small. We apply this basic idea in the proofs of our results.
In the following, we begin by introducing the core inequality in concentration of measure: Brunn-Minkowski. Then, we apply a variant of this inequality to prove Theorems 1 and 2. Followed by that, we present the preliminaries of Isoperimetric inequalities, and employ them to prove Theorem 3.
@@ 5.1 Brunn-Minkowski Inequality
For two subsets $A,B\subseteq\mathbb{R}^{d}$, their Minkowski sum is defined by $A+B=\left\{a+b:a\in A,b\in B\right\}$. The Brunn-Minkowski Inequality provides a crucial connection between volumes and Minkowski sum in Euclidean spaces.
Let $A$ and $B$ be two non-empty compact subsets of $\mathbb{R}^{d}$. The Brunn-Minkowski inequality claims that
@@
If one makes the additional assumption that $A$ and $B$ are restricted to convex subsets, then the inequality binds if and only if $A$ and $B$ are homothetic (that is one is the translated and scaled version of another). This inequality implies the concavity of the volume operator with respect to the Minkowski sum. There is a dimension-free version of this inequality that often proves more useful. In particular, for $\lambda\in[0,1]$, inequality (5.1) implies that
@@
We henceforth refer to this inequality by BM inequality. A useful application of BM is the following lemma: Expressing an upper bound for the minimum volume of two positively distanced subsets (its proof can be found in Artstein-Avidan et al. (2015), but we also state it for completeness).
@@
Since the volume of any Borel set can be approximated arbitrarily close by the inner measure of its closed subsets, we can assume without any loss that $A$ and $B$ are closed and hence compact. By the parallelogram law for the $\ell_{2}$-norm if $a\in A$ and $b\in B$ then
@@
where the inequality holds because $a,b\in\mathbb{B}$ and $\lVert a-b\rVert\geq\delta$. Hence, it follows that
@@
Setting $\lambda=1/2$ in (5.2) and using the above inequality justify the claim in (5.3). ∎
Therefore, as the dimension grows, two subsets in the $\ell_{2}$ ball with a bounded radius, will have positive distance from each other only if at least one them has a very small volume. Of course, the larger is the distance, the smaller would be the implied volume.
5.2 Proofs of the results in Section 3.1
We proceed with the proof of our first two theorems. In both cases, using the optimality or the equilibrium property of the allocation, we apply a type of convex separation argument. In the first theorem, the separation is provided by the given equilibrium price. While in the second theorem, the separation argument follows the idea used in the proof of the second welfare theorem. Consequently, as we argue below, the approximate versions of upper contour sets stay at a positive distance from their separating counterpart. An application of Lemma 1 on appropriately chosen subsets imply the volume bounds in Theorems 1 and 2.
@@
$\begin{gathered}H^{+}(p\,;\,b)\coloneqq\left\{x\in\mathbb{R}^{d}:p\,\cdot\,x\geq b\right\}\,,\\ H^{-}(p\,;\,b)\coloneqq\left\{x\in\mathbb{R}^{d}:p\,\cdot\,x\leq b\right\}\,,\end{gathered}$
that are, respectively, called upper and lower half-spaces. One can readily verify that the $\ell_{2}$ distance between the two half-spaces $H^{+}(p\,;\,b_{2})$ and $H^{-}(p\,;\,b_{1})$, where $b_{2}>b_{1}$, is equal to
@@
Proof of Theorem 1.
Since $f=\{f_{i}:i\in I\}$ is a Walrasian equilibrium allocation in $\mathcal{E}$, and preferences are monotone, there exists a price vector $p\in\mathbb{R}_{+}^{d}$ such that $p\,\cdot\,g_{i}>p\,\cdot\,\omega_{i}$ for all $i\in I$ and $g_{i}\in\mathcal{U}_{i}^{(0)}(f_{i})$. Next, observe that if $g\in\mathcal{U}_{i}^{(\varepsilon)}(f_{i})$ then $(1-\varepsilon)g\in\mathcal{U}_{i}^{(0)}(f_{i})$ and therefore $p\cdot\left((1-\varepsilon)(g-\omega_{i})-\varepsilon\omega_{i}\right)>0$. This in turn implies that
@@
where the last inequality holds because $p\in\mathbb{R}_{+}^{d}$ and $\omega_{i}\geq\tau\bm{1}$. Therefore, since $i\in I$ and $g\in\mathcal{U}_{i}^{(\varepsilon)}(f_{i})$ were arbitrary, by the above inequality $\mathcal{U}_{i}^{(\varepsilon)}(f_{i})-\{\omega_{i}\}\subseteq H^{+}\left(p;\varepsilon\tau\lVert p\rVert_{1}\right)$ for all $i\in I$. Let us define $\mathcal{Q}\coloneqq\bigcup_{i\in I}\left(\mathcal{U}_{i}^{(\varepsilon)}(f_{i})-\{\omega_{i}\}\right)$. [^9] The set $\mathcal{Q}$ was originally used by Debreu and Scarf (1963) to prove core convergence, and by Barman and Echenique (2023) to characterize approximate Walrasian equilibria. Then $\mathcal{Q}\subseteq H^{+}\left(p;\varepsilon\tau\lVert p\rVert_{1}\right)$, so for an arbitrary $r>0$, one has
@@
The equality above follows from (5.4), and the last inequality holds because $\min_{p\neq\bm{0}}\lVert p\rVert_{1}/\lVert p\rVert_{2}=1$, namely the minimum of $\lVert p\rVert_{1}/\lVert p\rVert_{2}$ is achieved on the standard unit basis vectors, and is equal to $1$. Now set $A\coloneqq\mathcal{Q}\cap\mathbb{B}(r)$ and $B\coloneqq H^{-}(p;0)\cap\mathbb{B}(r)$. By the above inequality $\mathsf{dist}(A,B)\geq\varepsilon\tau$. Since $p$ is a nonzero vector in $\mathbb{R}_{+}^{d}$, the subset $B$ covers at least half of the volume of $\mathbb{B}(r)$. So it must be that $\mathsf{Vol}(A)\leq\mathsf{Vol}(B)$. Therefore, Lemma 1 implies that $\mathsf{Vol}(A)/\mathsf{Vol}(\mathbb{B}(r))\leq\mathrm{e}^{-\varepsilon^{2}\tau^{2}d/8r^{2}}$, namely:
@@
Now observe that if $f=\{f_{i}:i\in I\}$ is a Walrasian equilibrium for the exchange economy $\mathcal{E}$, it is also a Walrasian equilibrium for the exchange economy $\mathcal{E^{\prime}}$ that is identical to $\mathcal{E}$ except that each agent $i$’s endowment is $\omega^{\prime}_{i}=f_{i}$. Therefore, we can replace $\omega_{i}$ in the above inequality with $f_{i}$, and obtain the volume bound in (3.2) that is equivalent to (3.1).∎
Proof of Corollary 1.
The union bound and (3.1) yields

### v4

@@ 3.3 Prior Beliefs and Welfare-Improving Trade
Our third result concerns agents with multiple priors, and the size of the sets of prior beliefs that they may possess. In previous sections, we quantified the upper contour sets of individual agents and their sums. Here we instead follow Yaari (1969) to interpret a vector that supports the upper contour sets, at some contingent consumption, as a prior belief. We depend on a relaxation of the equivalence between the existence of a common prior and Pareto efficiency of an allocation. This equivalence has been studied in a number of previous works (e.g., Billot et al., 2000; Ng, 2003; Rigotti et al., 2008; Gilboa et al., 2014; Ghirardato and Siniscalchi, 2018). We shall follow Rigotti et al. (2008) quite closely here.
We consider an exchange economy $\mathcal{E}$ with no aggregate uncertainty. Importantly, from this point on we will not require convexity of preferences. The aggregate endowment is the same across all states of the world, i.e., $\omega=(\bar{\omega},\ldots,\bar{\omega})$. Also following the convention of Rigotti et al. (2008), in this section, we restrict the allocation space to nonnegative vectors. In this sense, we denote the unrestricted allocation space (defined in Section 2.3) by $\mathcal{F}_{\bar{\omega}}$, and its nonnegative subset by $\mathcal{F}^{+}_{\bar{\omega}}\coloneq\mathcal{F}_{\bar{\omega}}\cap\mathbb{R}_{+}^{d}$. We quantify its “size” by
$\rho\coloneq 2\bar{\omega}^{-1}\max_{f\in\mathcal{F}^{+}_{\bar{\omega}}}\sum_{i\in I}\lVert f_{i}\rVert\,.$ (3.7)
@@
The set $\mathcal{B}_{i}(f)$ is a convex and closed (hence compact) subset of the $(d-1)$-dimensional probability simplex $\Delta S$ (or $\Delta_{d}$). One may also interpret a vector in $\mathcal{B}_{i}(f)$ as the vector of normalized prices that supports the consumption of the act $f$.
@@
$\mathcal{B}_{i}(f_{i})^{\delta}\coloneq\{\nu\in\Delta S:\inf_{\mu\in\mathcal{B}_{i}(f_{i})}\lVert\nu-\mu\rVert<\delta\}\,.$
@@
Leveraging Proposition 2, in the following theorem we examine the volume of the prior sets as the number of states $d$ grows large. Before we state the result, we need to introduce some further notations. For a subset $J\subseteq I$, denote its complement by $J^{c}$, and define $\mathcal{B}_{J}(f_{J})\coloneq\bigcap_{j\in J}\mathcal{B}_{j}(f_{j})$. We often drop the allocation $f$ from the argument of subjective belief sets, when it is understood from the context. We further assume that preferences are strongly monotonic (i.e., belonging to $\mathcal{P}^{\text{sm}}$). This assumption ensures that the dimension of each $\mathcal{B}_{i}$ is the same as $\Delta_{d}$.
@@
Let $\mathcal{E}\colon I\to\mathcal{P}^{\text{sm}}\times\mathbb{R}_{+}^{d}$ be an exchange economy with no aggregate uncertainty. If the allocation $f\in\mathcal{F}^{+}_{\bar{\omega}}$ is $\varepsilon$-Pareto dominated, then there exists a constant $c>0$, such that for every proper subset $J\subset I$,
@@
The message of the theorem is clearest in a two-agent economy (for example the setting of Kajii and Ui (2006) with MEU preferences). The existence of an $\varepsilon$-Pareto improving trade is then possible only when one of the two agents has a small set of subjective beliefs, as measured through its Euclidean volume. We expand on this interpretation in Section 4.1.
More generally, in an economy with many agents, Theorem 3 states that, regardless of how we partition the agents into two groups, at least one group must have a small set of priors—meaning the volume of the intersection of their prior sets is small. If we think of the intersection as corresponding to a representative agent for the group (a notion that may be formalized), then this agent must be close to being ambiguity neutral: see again the discussion in Section 4.1 for a sufficient condition under which the small volume conclusion leads to a behavioral property that bounds the degree of ambiguity aversion.
@@ 4.1 Behavioral Implications of Theorem 3
In this section we study the behavioral implications of Theorem 3. For the purposes of this discussion, we consider a setting with two agents. The message of our theorem is that, if the current risk-sharing arrangement between the two agents is strongly Pareto dominated (namely if there exists a possibility of a strongly welfare improving ex-ante trade), then the volume of the underlying belief set of at least one agent must be very small. The smaller this volume is, the closer that agent is to being ambiguity neutral—a connection we seek to quantify in this section. [^8] For an overview of ambiguity and ambiguity aversion, see Machina and Siniscalchi (2014) and Gilboa and Marinacci (2016).
@@
We further assume that $\Pi$ has constant width, namely that the distance between any distinct parallel supporting hyperplanes of $\Pi$ (residing on the $(d-1)$-dimensional probability simplex) is constant. This assumption is crucial for our subsequent use of the result of Schramm (1988), but one can imagine a quantitative relaxation of the assumption of constant width and that it would translate into a modified bound on the degree of ambiguity aversion.
@@
$\theta(f)\coloneq\max_{\mu\in\Pi}f\cdot\mu-\min_{\mu\in\Pi}f\cdot\mu\,.$
When $\Pi$ has constant width, $\theta(f)$ becomes constant in $f$, and we write $\theta(f)\equiv\theta$. Thus we can take $\theta$ as a measure of ambiguity aversion. Its meaning is the difference in how an ambiguity averse and ambiguity loving agent with the same set of priors would evaluate any given normalized act. A small degree of ambiguity aversion corresponds to a small difference between the ambiguity averse and ambiguity loving evaluations of an act.
@@
Under the conditions of Theorem 3, let $\Pi$ coincide with the set of priors with the smaller volume, and suppose that it has a constant width $\theta$. Then, there exists a universal constant $c>0$ such that
@@
4.2 On the Interpretation and Magnitude of $\varepsilon$
The presence of $\varepsilon$ in a welfare improvement can be justified by appealing to robustness or transaction costs. When we talk about renegotiation, it is natural to imagine that reaching an agreement, with the requisite (unmodeled) communication requirements, is costly and is only worthwhile if the agents can obtain an improvement by some given margin. The number $\varepsilon>0$ captures such transaction costs. Alternatively, we may restrict attention to some robust optimization criterion, which demands a minimum utility gain before moving away from a status quo.
The interpretation of our results depends on the magnitude of two numbers: the number of states $d$, and the fraction $\varepsilon$ accounted for in a utility improvement. When $d$ is large while $\varepsilon$ remains constant, the probabilities discussed in Theorems 1 and 2, and the volume in Theorem 3, shrink to zero. So one could question our interpretation by arguing that we should use a small value of $\varepsilon$ when the number of states $d$ is large. In particular, one might argue that we should impose that $\varepsilon=O(1/\sqrt{d})$. We would disagree, however, essentially because we think of $\varepsilon$ as a dimension-free fraction.
@@ 4.2 On the Interpretation and Magnitude of $\varepsilon$
Second, taking $\varepsilon=O(1/\sqrt{d})$ is problematic because it seems very hard to reconcile with the common practice of using a numerical objective function in calculating approximately optimal outcomes. As we discussed above, many applications make use of a homothetic preference and a resulting utility function that is homogeneous. Examples include the max-min representation in choice under uncertainty, or the Cobb-Douglas utility in consumer choice. In this case, the $\varepsilon$ equals that tolerance level assumed in the agents’ maximization problem: the number $\varepsilon$ is then measured in “utils,” the same unit of account used for the utility function. Utils are, however, dimension-free: for example, if $u$ is a Cobb-Douglas utility, then $u(\mathbf{1})=1$ regardless of $d$. In a model with many states of the world, we would hardly be allowing for any relaxation in our notion of approximate optimality.
Finally, we should emphasize the reinterpretation of our results using the coefficient of resource utilization (see Proposition 1). The CRU is usually thought of as a fraction of national income, also a dimension-free measure, and one that might be expected to be constant, or even grow, with the size of the economy.
@@ 5 Methodology and Proofs
At a high level, the primary concept underpinning our results is the concentration of measure phenomenon. Every economist is familiar with the idea that, with a large sample, some sample statistics (such as the mean) are with high probability close to their population counterparts. In Euclidean spaces, dimension can play the role of sample size: in high-dimensional Euclidean spaces, and under relatively mild conditions, probability measures tend to concentrate as well. Specifically, for a probability measure $\mu$ on a bounded set $K\subseteq\mathbb{R}^{d}$, and a subset $A$ that encompasses at least half of the probability space, namely $\mu(A)\geq 1/2$, the metric extension $A^{\delta}=\{z:\mathsf{dist}(z,A)<\delta\}$ covers a substantial and growing portion of the measure $\mu(K)$. In particular, the complement $1-\mu(A^{\delta})$ diminishes rapidly, often exhibiting exponential decay with respect to the dimension $d$. These types of concentration bounds are commonly referred to as isoperimetric inequalities.
The importance of high-dimensional isoperimetric inequalities for our work lies in the independence of the concentration rate from the subset $A$. Using the words we introduced earlier in the paper: shapes do not matter. The source of our results—that certain welfare improvements have a vanishingly small probability, independent of the shape of the agents’ preferences—can be traced to basic observations in the theory of isoperimetric inequalities and concentration of measures.
A consequence of the geometric concentration of measure is that, if two subsets $A$ and $B$ are separated with a positive distance, then as the dimension $d$ grows, the measure of at least one of them must be exponentially small. We apply this idea in the proofs of our results.
In this section, we begin by introducing the core inequality that relates to the concentration of measures: the Brunn-Minkowski inequality. Then, we apply a variant of this inequality to prove Theorems 1 and 2. Once we have shown these results, we present the preliminaries of isoperimetric inequalities, and employ them to prove Theorem 3.
@@ 5.1 Brunn-Minkowski Inequality
For two subsets $A,B\subseteq\mathbb{R}^{d}$, their Minkowski sum is defined by $A+B\coloneq\left\{a+b:a\in A,b\in B\right\}$. The Brunn-Minkowski inequality provides a crucial connection between volumes and the Minkowski sum in Euclidean spaces.
Let $A$ and $B$ be two non-empty compact subsets of $\mathbb{R}^{d}$. The Brunn-Minkowski inequality (Gardner, 2002) claims that
@@
If one makes the additional assumption that $A$ and $B$ are restricted to convex subsets, then the inequality binds if and only if $A$ and $B$ are homothetic (that is, one is the translated and scaled version of the other). This inequality implies the concavity of the volume operator with respect to the Minkowski sum. There is a dimension-free version of this inequality that often proves more useful. In particular, for $\lambda\in[0,1]$, inequality (5.1) implies that
@@
We henceforth refer to this inequality as the BM inequality. A useful application of the BM inequality is the following lemma, which provides an upper bound for the minimum volume of two positively distanced subsets (its proof can be found in Artstein-Avidan et al. (2015), but we include it here for completeness).
@@
Since the volume of any Borel set can be approximated arbitrarily closely by the inner measure of its closed subsets, we can assume without loss of generality that $A$ and $B$ are closed, and hence compact. By the parallelogram law for the $\ell_{2}$-norm if $a\in A$ and $b\in B$ then
@@
where the inequality holds because $a,b\in\mathbb{B}(r)$ and $\lVert a-b\rVert\geq\delta$. Hence, it follows that
@@
Setting $\lambda=1/2$ in (5.2) and using the above inequality justifies the claim in (5.3). ∎
Therefore, as the dimension grows, two subsets in the $\ell_{2}$ ball with a bounded radius will have positive distance from each other only if at least one of them has a very small volume. Of course, the greater the distance, the smaller the implied volume.
5.2 Proof of the Results in Sections 3.1 and 3.2
We proceed with the proof of our first two theorems. In both cases, using the optimality or the equilibrium property of the allocation, we apply a type of convex separation argument. In the first theorem, the separation is provided by the given equilibrium price. While in the second theorem, the separation argument follows the idea used in the proof of the second welfare theorem. Consequently, as we argue below, the approximate versions of upper contour sets stay at a positive distance from their separating counterpart. An application of Lemma 1 to appropriately chosen subsets implies the probability bounds in Theorems 1 and 2.
@@
$\begin{gathered}H^{+}(p\,;\,b)\coloneq\left\{x\in\mathbb{R}^{d}:p\,\cdot\,x\geq b\right\}\,,\\ H^{-}(p\,;\,b)\coloneq\left\{x\in\mathbb{R}^{d}:p\,\cdot\,x\leq b\right\}\,,\end{gathered}$
that are, respectively, called upper and lower half-spaces. One can readily verify that the $\ell_{2}$ distance between the two parallel half-spaces $H^{+}(p\,;\,b_{2})$ and $H^{-}(p\,;\,b_{1})$, where $b_{2}>b_{1}$, is equal to
@@
Proof of Theorem 1
Since $f=(f_{i})_{i\in I}$ is a Walrasian equilibrium allocation in $\mathcal{E}$, and preferences are monotone, there exists a price vector $p\in\mathbb{R}_{+}^{d}$ such that $p\,\cdot\,g_{i}>p\,\cdot\,\omega_{i}$ for all $i\in I$ and $g_{i}\in\mathcal{U}_{i}^{(0)}(f_{i})$. Next, observe that if $g\in\mathcal{U}_{i}^{(\varepsilon)}(f_{i})$ then $(1-\varepsilon)g\in\mathcal{U}_{i}^{(0)}(f_{i})$ and therefore $p\cdot\left((1-\varepsilon)(g-\omega_{i})-\varepsilon\omega_{i}\right)>0$. This in turn implies that
@@
where the last inequality holds because $p\in\mathbb{R}_{+}^{d}$ and $\omega_{i}\geq\tau\bm{1}$. Therefore, since $i\in I$ and $g_{i}\in\mathcal{U}_{i}^{(\varepsilon)}(f_{i})$ were arbitrary, by the above inequality $\mathcal{U}_{i}^{(\varepsilon)}(f_{i})-\{\omega_{i}\}\subseteq H^{+}\left(p;\varepsilon\tau\lVert p\rVert_{1}\right)$ for all $i\in I$. Let us define $\mathcal{Q}\coloneq\bigcup_{i\in I}\left(\mathcal{U}_{i}^{(\varepsilon)}(f_{i})-\{\omega_{i}\}\right)$. [^9] The set $\mathcal{Q}$ was originally used by Debreu and Scarf (1963) to prove core convergence, and by Barman and Echenique (2023) to characterize approximate Walrasian equilibria. Then $\mathcal{Q}\subseteq H^{+}\left(p;\varepsilon\tau\lVert p\rVert_{1}\right)$, so for an arbitrary $r>0$, one has
@@
The equality above follows from (5.4), and the last inequality holds because $\min_{p\neq\bm{0}}\lVert p\rVert_{1}/\lVert p\rVert_{2}=1$, namely the minimum of $\lVert p\rVert_{1}/\lVert p\rVert_{2}$ is achieved on the standard unit basis vectors, and is equal to $1$. Now set $A\coloneq\mathcal{Q}\cap\mathbb{B}(r)$ and $B\coloneq H^{-}(p;0)\cap\mathbb{B}(r)$. By the above inequality $\mathsf{dist}(A,B)\geq\varepsilon\tau$. Since $p$ is a nonzero vector in $\mathbb{R}_{+}^{d}$, the subset $B$ covers at least half of the volume of $\mathbb{B}(r)$. So it must be that $\mathsf{Vol}(A)\leq\mathsf{Vol}(B)$. Therefore, Lemma 1 implies that $\mathsf{Vol}(A)/\mathsf{Vol}(\mathbb{B}(r))\leq\mathrm{e}^{-\varepsilon^{2}\tau^{2}d/8r^{2}}$, namely:
@@
Now observe that if $f=(f_{i})_{i\in I}$ is a Walrasian equilibrium for the exchange economy $\mathcal{E}$, it is also a Walrasian equilibrium for the exchange economy $\mathcal{E^{\prime}}$ that is identical to $\mathcal{E}$ except that each agent $i$’s endowment is $\omega^{\prime}_{i}=f_{i}$. Therefore, we can replace $\omega_{i}$ in the above inequality with $f_{i}$, and obtain the volume bound in (3.2) that is equivalent to
$\mathsf{P}^{r}\left((1-\varepsilon)(f_{i}+z)\succ_{i}f_{i}\text{ for some }i\in I\right)\leq\mathrm{e}^{-\varepsilon^{2}\tau^{2}d/8r^{2}}\,.$
Since $\frac{\mathrm{d}\mathsf{P}^{r}_{\kappa}}{\mathrm{d}\mathsf{P}^{r}}(z)\leq\kappa$ almost everywhere on $z\in\mathbb{B}(r)$, the claim in (3.1) follows from the above bound.∎
Proof of Corollary 1
The union bound and (3.2) imply that